# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v2

@@ 3.1 Entity Alignment Model
The entity alignment (EA) model aims to learn informative entity embeddings and perform model training for entity alignment inference. Our UPL-EA framework is designed to be modular, allowing for the use of various existing EA models. In this work, we present one instantiation based on our previous work [10].
3.1.1 Entity Embedding
To enable the EA model to learn informative entity embeddings, it is essential to capture structural information inherent in KGs. In this work, we employ a global-local neighborhood aggregation scheme [10] that has the advantages of encoding both global and local patterns in KG structures.
Global-Level Relation Aggregation. To capture the global patterns of a given relation $r$, we construct its feature vector $\bm{x}_{r}$ by considering all triplets associated with $r$ in two KGs. Specifially, for each relation $r\in\mathcal{R}_{1}\cup\mathcal{R}_{2}$, we compute its feature vector $\bm{x}_{r}$ by averaging the concatenated feature vectors of its corresponding head and tail entities:
$\bm{x}_{r}=\text{Average}\bigl{(}\{[\bm{x}_{i}\|\bm{x}_{j}]|\;(e_{i},r,e_{j})\in\mathcal{T}_{1}\cup\mathcal{T}_{2}\}\bigl{)},$ (1)
[…]
$\bm{x}_{i}^{\text{rel}}=\text{Average}\bigl{(}\{\bm{x}_{r}|\;r\in\mathcal{N}_{r}(e_{i})\}\bigl{)},$ (2)
where $\mathcal{N}_{r}(e_{i})$ is the set of one-hop neighboring relations of entity $e_{i}$.
To perform global-level relation aggregation, we concatenate each entity’s averaged neighboring relation feature $\bm{x}_{i}^{\text{rel}}\in\mathbb{R}^{2f}$ with its original feature $\bm{x}_{i}\in\mathbb{R}^{f}$, followed by a non-linear transformation. To avoid over-smoothing, the original entity feature vector $\bm{x}_{i}\in\mathbb{R}^{f}$ is also added via a residual connection. The relation aggregated entity embedding is obtained as:
$\bm{h}_{i}^{(1)}=\mathrm{ReLU}\bigl{(}\bm{W}[\bm{x}_{i}\|\bm{x}_{i}^{\text{rel}}]+\bm{b}\bigl{)}+\bm{x}_{i},$ (3)
where $\bm{h}_{i}^{(1)}\in\mathbb{R}^{d}$ and $d$ is the dimension of entity embedding. $\bm{W}\in\mathbb{R}^{d\times 3f}$ and $\bm{b}\in\mathbb{R}^{d}$ are learnable parameters. Generally, the dimensions of entity embedding and entity feature are different. Here, for the residual connection, we set $d=f$.
Local-Level Entity Aggregation. After obtaining entity embeddings via relation aggregation at the global level, we further perform local-level entity aggregation to capture neighboring entity structures. For this purpose, we employ a two-layer GCN connected by a highway gate strategy [34] to avoid over-smoothing. Formally, the entity embedding vector $\bm{h}_{i}^{(1)}$ is updated as follows:
$\bm{h}_{i}^{(l+1)}=\text{Aggregate}\bigl{(}\{\bm{h}_{j}^{(l)}|\;e_{j}\in\mathcal{N}_{e}(e_{i})\cup\{e_{i}\}\}\bigl{)},$ (4)
where $\mathcal{N}_{e}(e_{i})$ is the set of one-hop neighboring entities of entity $e_{i}$, $\text{Aggregate}(\cdot)$ denotes the aggregation function specific to a high-way GCN encoder. Finally, the embedding for entity $e_{i}$ is obtained as $\bm{h}_{i}=\bm{h}_{i}^{(3)}\in\mathbb{R}^{d}$.
3.1.2 Model Training for Embedding Learning
Given a set of seed alignments $\mathcal{S}$, we define a margin-based loss function for embedding learning, such that the equivalent entities are encouraged to be close to each other in the embedding space:
@@ 3.1.2 Model Training for Embedding Learning
where $\mathcal{S}^{-}_{(e_{i},e_{j})}=\{(e_{i},e_{j}^{-})|\;e_{j}^{-}\in\mathcal{E}_{2}\setminus\{e_{j}\}\}\cup\{(e_{i}^{-},e_{j})|\;e_{i}^{-}\in\mathcal{E}_{1}\setminus\{e_{i}\}\}$ denotes the set of negative alignments, synthesized by negative sampling of a positive alignment $(e_{i},e_{j})\in\mathcal{S}$ as $(e_{i},e_{j}^{-})\notin\mathcal{S}$ and $(e_{i}^{-},e_{j})\notin\mathcal{S}$. $\gamma$ is a hyper-parameter that determines the margin to separate positive alignments from negative alignments. $d(e_{i},e_{j})$ indicates the embedding distance between entity pair $(e_{i},e_{j})$ across two KGs, defined as:
@@
The loss function in Eq. (5) can be minimized with respect to entity embeddings $\{\bm{h}_{i}|\;e_{i}\in\mathcal{E}_{1}\cup\mathcal{E}_{2}\}$. To facilitate model training, we adopt an adaptive negative sampling strategy to obtain the set of negative alignments $\mathcal{S}^{-}_{(e_{i},e_{j})}$. Specifically, for each positive alignment $(e_{i},e_{j})\in\mathcal{S}$, we select $K$ nearest entities of $e_{i}$ (or $e_{j}$), measured using the embedding distance in Eq. (6), to replace $e_{j}$ (or $e_{i}$) and form $K$ negative counterparts $(e_{i},e_{j}^{-})$ (or $(e_{i}^{-},e_{j})$). This adaptive strategy helps generate “hard” negative alignments and push their associated entities to be apart from each other in the embedding space.
@@ 3.2 Unified Pseudo-Labeling Strategy
After training the EA model, the learned entity embeddings can be used for alignment inference. However, as prior seed alignments provided for training are often limited, the performance of entity alignment can be suboptimal. Therefore, pseudo-labeling strategies can be designed to select a set of unaligned entity pairs as pseudo-labeled alignments, which are used to augment prior seed alignments for boost model training. However, as discussed in Section 2.2, a naive strategy might introduce a large number of pseudo-labeling errors, leading to confirmation bias. To address this, we propose the Unified Pseudo-Labeling (UPL) that explicitly aims to mitigate Type I and Type II pseudo-labeling errors.
In what follows, the two essential components of UPL: OT-based pseudo-labeling and pseudo-label ensembling, are discussed in detail.
@@ 3.2.1 OT-Based Pseudo-Labeling
To mitigate Type I pseudo-labeling errors, we propose using Optimal Transport (OT) as an effective means to reliably pseudo-label unaligned entities across KGs to reinforce the training of the EA model. As a powerful mathematical framework for transforming one distribution into another, OT allows us to more effectively identify one-to-one correspondences between entities, ensuring a coherent alignment configuration.
Formally, we model the alignment between two unaligned entity sets $\mathcal{E}^{U}_{1}$ and $\mathcal{E}^{U}_{2}$ as an OT process to warrant one-to-one alignments, i.e., transporting each entity $e_{i}\in\mathcal{E}^{U}_{1}$ to a unique entity $e_{j}\in\mathcal{E}^{U}_{2}$, with minimal overall transport cost. Denote $\bm{C}\in\mathbb{R}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ as the transport cost matrix, and without loss of generality, we assume $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$. The transport plan is a mapping function $T:e_{i}\rightarrow T(e_{j})$, where $e_{i}\in\mathcal{E}^{U}_{1}$, $T(e_{i})\in\mathcal{E}^{U}_{2}$. Thus, the objective of entity alignment is to find the optimal transport plan $T^{*}$ that minimizes the overall transport cost:
@@
A critical aspect of the above objective is defining a reliable measure of the transport cost. One might directly use the distances between the learned entity embeddings. However, when only a limited set of prior seed alignments are available for training, the learned entity embeddings can be uninformative, particularly during the early training stages before the EA model has converged. As a result, using these embeddings to calculate the distances for defining the transport cost can be error-prone. To address this, we resort to rectifying the embedding distance using relational neighborhood matching [62], which complements the training of the EA model, especially during the early stages, for learning better entity embeddings and providing a more reliable cost measure for OT modeling. The principle of distance rectification is to leverage relational contexts within local neighborhoods to help determine the extent to which two entities should be aligned. Intuitively, if two entities $e_{i}\in\mathcal{E}_{1}$ and $e_{j}\in\mathcal{E}_{2}$ share more aligned neighboring entities/relations, the distance between their embeddings should be smaller, indicating a higher likelihood of being aligned to each other. Based on this intuition, the transport cost for OT-based pseudo-labeling is defined as follows:
@@
To solve the discrete OT problem in Eq. (10), several exact algorithms have been proposed, such as interior point methods [49] and network simplex [25]. While these exact algorithms guarantee to find a closed-form optimal transport plan, their high computational cost makes them intractable for iterative pseudo-labeling. Thus, we propose to use an entropy regularized OT problem, as defined in Eq. (11) below, which can be solved by the efficient Sinkhorn algorithm  [8].
@@
Once the optimal coupling matrix $\bm{P}^{*}$ is estimated, entity alignments can be inferred accordingly. Since one-to-one correspondences are crucial for eliminating conflicted misalignments (Type I pseudo-labeling errors), we further propose a selection criterion to identify entity pairs as pseudo-labeled alignments:
@@
Since the coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ has non-negative entries, the summation $\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$ from Eq. (3.2.1) must be no smaller than any component of it, i.e., $\bm{P}^{*}_{e_{i},e_{k}}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. Together with Eq. (3.2.1), we can further derive that any component in the summation is smaller than the decision threshold, i.e., $\bm{P}^{*}_{e_{i},e_{k}}{\color[rgb]{0,0,0}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. In other words, all other probability values in the same row of $\bm{P}^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold. Thus, no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments.
@@
The OT-based pseudo-labeling algorithm is provided in Algorithm 1. In Step 1, the algorithm starts by calculating the transport cost matrix $\bm{C}$, with a time complexity of $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$, where $d$ is the embedding dimension. In Steps 2-9, the Sinkhorn algorithm takes the transport cost matrix $\bm{C}$ as input to estimate the optimal transport plan $\bm{P}^{*}$ via iterative row normalization and column normalization, the time complexity is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|/\beta)$. In Step 10, entity pairs with values in $\bm{P}^{*}$ larger than the decision threshold are selected as pseudo-labeled alignments. Finally, in Step 11, the algorithm returns a set of pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$. The overall time complexity of Algorithm 1 is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$.
@@ 7       $\bm{P}=\bm{a}\bm{b}^{\top}\odot\bm{Q}$;
After obtaining the set of pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$, seed alignments are updated as $\mathcal{S}\leftarrow\mathcal{S}_{0}\cup\widehat{\mathcal{S}}_{t}$ for training the EA model in the next iteration. To distinguish the contributions of pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$ from the prior seed alignments $\mathcal{S}_{0}$ provided beforehand during model training, we assign a soft weighting score to each alignment provided for training, defined as follows:
$w(e_{i},e_{j})=\begin{cases}1,&(e_{i},e_{j})\in\mathcal{S}_{0},\\ 1/(1+e^{\bm{C}_{e_{i},e_{j}}}),&(e_{i},e_{j})\in\widehat{\mathcal{S}}_{t}.\\ \end{cases}$ (14)
For each pseudo-labeled alignment $(e_{i},e_{j})\in\widehat{\mathcal{S}}_{t}$, its weighting score $w(e_{i},e_{j})$ is within the range of $(0,1)$ and inversely related to the transport cost $\bm{C}_{e_{i},e_{j}}$. This ensures that the contribution of each pseudo-labeled alignment is appropriately moderated compared to the ground truth alignments during model training. Accordingly, the margin-based loss function for embedding learning, defined in Eq. (5), is updated as follows:
$L=\sum_{(e_{i},e_{j})\in\mathcal{S}}\sum_{(e_{i}^{-},e_{j}^{-})\in\mathcal{S}^{-}_{(e_{i},e_{j})}}w(e_{i},e_{j})\cdot\max\bigl{(}0,d(e_{i},e_{j})-d(e_{i}^{-},e_{j}^{-})+\gamma\bigl{)}.$ (15)
3.2.2 Pseudo-Label Ensembling
Through OT-based pseudo-labeling, a set of conflict-free pseudo-labeled alignments are inferred at each iteration. However, they are still susceptible to one-to-one misalignments (Type II pseudo-labeling errors), especially when the learned entity embeddings are uninformative at early training stages. To address this issue, we propose a self-ensembling approach that refines pseudo-labeled alignments by combining predictions over multiple pseudo-labeling iterations. The objective is to seek consensus predictions over multiple iterations to reduce the prediction variability, thereby improving the quality of pseudo-labeled alignments.

### v3

@@ 3.1 Entity Alignment Model
The entity alignment (EA) model aims to learn informative entity embeddings and perform model training for entity alignment inference.
To enable the EA model to learn informative entity embeddings, our UPL-EA framework adopts a modular entity embedding encoder, $f_{\textrm{en}}$, designed to effectively capture structural information inherent in KGs. Given an entity $e_{i}\in\mathcal{E}_{1}\cup\mathcal{E}_{2}$ with its associated feature vector $\bm{x}_{i}\in\mathbb{R}^{n}$, the encoder $f_{\textrm{en}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ maps the entity into a $d$-dimensional embedding space $\mathbb{R}^{d}$:
$\bm{h}_{i}=f_{\textrm{en}}(\bm{x}_{i},\mathcal{G}_{1},\mathcal{G}_{2};\bm{\Theta}),$ (1)
where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are the two KGs, $\bm{\Theta}$ represents the learnable parameters of the EA model, and $\bm{h}_{i}\in\mathbb{R}^{d}$ is the encoded entity embedding. The encoder $f_{\textrm{en}}$ can be instantiated with any expressive entity embedding model. In this work, we adopt a highway-gated GCN with a global-local neighborhood aggregation scheme, following our previous work [10].
After obtaining the encoded entity embeddings $\{\bm{h}_{i}|\;e_{i}\in\mathcal{E}_{1}\cup\mathcal{E}_{2}\}$, we then define a margin-based loss function for embedding learning, such that the equivalent entities are encouraged to be close to each other in the embedding space:
@@
where $\mathcal{S}$ denotes a set of prior seed alignments initially provided for training. $\mathcal{S}^{-}_{(e_{i},e_{j})}=\{(e_{i},e_{j}^{-})|\;e_{j}^{-}\in\mathcal{E}_{2}\setminus\{e_{j}\}\}\cup\{(e_{i}^{-},e_{j})|\;e_{i}^{-}\in\mathcal{E}_{1}\setminus\{e_{i}\}\}$ denotes the set of negative alignments, synthesized by negative sampling of a positive alignment $(e_{i},e_{j})\in\mathcal{S}$ as $(e_{i},e_{j}^{-})\notin\mathcal{S}$ and $(e_{i}^{-},e_{j})\notin\mathcal{S}$. $\gamma$ is a hyper-parameter that determines the margin that separates positive alignments from negative alignments. $d(e_{i},e_{j})$ indicates the embedding distance between entity pair $(e_{i},e_{j})$ across two KGs, defined as:
@@
The loss function in Eq. (2) can be minimized with respect to entity embeddings $\{\bm{h}_{i}|\;e_{i}\in\mathcal{E}_{1}\cup\mathcal{E}_{2}\}$. To facilitate model training, we adopt an adaptive negative sampling strategy to obtain a set of negative alignments $\mathcal{S}^{-}_{(e_{i},e_{j})}$. Specifically, for each positive alignment $(e_{i},e_{j})\in\mathcal{S}$, we select $K$ nearest entities of $e_{i}$ (or $e_{j}$), measured using the embedding distance in Eq. (3), to replace $e_{j}$ (or $e_{i}$) and form $K$ negative counterparts $(e_{i},e_{j}^{-})$ (or $(e_{i}^{-},e_{j})$). This strategy helps generate “hard” negative alignments and pushes their associated entities to be apart from each other in the embedding space.
@@ 3.2 Unified Pseudo-Labeling Strategy
After training the EA model, the learned entity embeddings can be used for alignment inference. However, as prior seed alignments initially provided for training are often limited, the performance of entity alignment can be suboptimal. Therefore, pseudo-labeling strategies can be designed to select a set of unaligned entity pairs as pseudo-labeled alignments, which are used to augment seed alignments for boosting model training. However, as previously discussed in Section 2.2, a naive strategy inevitably introduces a considerable number of pseudo-labeling errors, leading to confirmation bias. To address this issue, we propose UPL, a unified pseudo-labeling framework that explicitly aims to mitigate conflicted misalignments and one-to-one misalignments.
In what follows, the two key components of UPL: OT-based pseudo-labeling and parallel pseudo-label ensembling, are discussed in detail.
@@ 3.2.1 OT-based Pseudo-Labeling
To mitigate conflicted misalignments, we propose using Optimal Transport (OT) as an effective means to reliably pseudo-label unaligned entity pairs across KGs to reinforce the training of the EA model. As a powerful mathematical framework for transforming one distribution into another, OT allows to more effectively identify one-to-one correspondences between entities, ensuring a coherent alignment configuration.
Formally, we model the alignment between two unaligned entity sets $\mathcal{E}^{U}_{1}$ and $\mathcal{E}^{U}_{2}$ as an OT process to warrant the one-to-one correspondence, i.e., transporting each entity $e_{i}\in\mathcal{E}^{U}_{1}$ to a unique entity $e_{j}\in\mathcal{E}^{U}_{2}$, with minimal overall transport cost. Denote $\bm{C}\in\mathbb{R}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ as the transport cost matrix, and without loss of generality, we assume $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$. The transport plan is a mapping function $T:e_{i}\rightarrow T(e_{j})$, where $e_{i}\in\mathcal{E}^{U}_{1}$, $T(e_{i})\in\mathcal{E}^{U}_{2}$. Thus, the objective of entity alignment is to find the optimal transport plan $T^{*}$ that minimizes the overall transport cost:
@@
A critical aspect of the above objective is defining a reliable measure of the transport cost. One might directly use the distances between the learned entity embeddings. However, when only a limited number of prior seed alignments are available for training, the learned entity embeddings can be uninformative, particularly during the early training stages before the EA model has converged. As a result, using these embeddings to calculate the distances for defining the transport cost can be error-prone. To address this, we resort to rectifying the embedding distance using relational neighborhood matching [63], which complements the training of the EA model, especially during the early stages, for learning better entity embeddings and providing a more reliable cost measure for OT modeling. The principle of distance rectification is to leverage relational contexts within local neighborhoods to help determine the extent to which two entities should be aligned. Intuitively, if two entities $e_{i}\in\mathcal{E}_{1}$ and $e_{j}\in\mathcal{E}_{2}$ share more aligned neighboring entities/relations, the distance between their embeddings should be smaller, indicating a higher likelihood of being aligned to each other. Based on this intuition, the transport cost for OT-based pseudo-labeling is defined as follows:
@@
To solve the discrete OT problem in Eq. (7), several exact algorithms have been proposed, such as interior point methods [50] and network simplex [26]. While these exact algorithms guarantee to find a closed-form optimal transport plan, their high computational cost makes them intractable for iterative pseudo-labeling. Thus, we propose to use an entropy regularized OT problem, as defined in Eq. (8) below, which can be solved by the efficient Sinkhorn algorithm [8]:
@@
Once the optimal coupling matrix $\bm{P}^{*}$ is estimated, entity alignments can be inferred accordingly. Since one-to-one correspondences are crucial for eliminating conflicted misalignments, we further propose a selection criterion to identify entity pairs as pseudo-labeled alignments:
@@
Since the coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ has non-negative entries, the summation $\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$ from Eq. (3.2.1) must be no smaller than any of its components, i.e., $\bm{P}^{*}_{e_{i},e_{k}}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. Please note that, as we focus on a transductive semi-supervised setting, the size of two unaligned entity sets, $|\mathcal{E}^{U}_{1}|$ and $|\mathcal{E}^{U}_{2}|$, are known. Together with Eq. (3.2.1), we can further derive that any component in the summation is smaller than the decision threshold, i.e., $\bm{P}^{*}_{e_{i},e_{k}}{\color[rgb]{0,0,0}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. In other words, all other probability values in the same row of $\bm{P}^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold. Thus, no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments.
@@
The OT-based pseudo-labeling algorithm is provided in Algorithm 1. In Step 1, the algorithm starts by calculating the transport cost matrix $\bm{C}$, with a time complexity of $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$, where $d$ is the embedding dimension. In Steps 2-9, the Sinkhorn algorithm takes the transport cost matrix $\bm{C}$ as input to estimate the optimal transport plan $\bm{P}^{*}$ via iterative row normalization and column normalization, the time complexity is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|/\beta)$. In Step 10, entity pairs with values in $\bm{P}^{*}$ larger than the decision threshold are selected as pseudo-labeled alignments. Finally, in Step 11, the algorithm returns a set of conflict-free pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$. The overall time complexity of Algorithm 1 is $O(|\mathcal{E}^{U}_{1}|\cdot|\mathcal{E}^{U}_{2}|\cdot d)$.
@@ 7       $\bm{P}=\bm{a}\bm{b}^{\top}\odot\bm{Q}$;
3.2.2 Parallel Pseudo-Label Ensembling
Through OT-based pseudo-labeling, conflict-free pseudo-labeled alignments are inferred. However, these alignments remain susceptible to one-to-one misalignments, particularly when the learned entity embeddings are still uninformative in the early stages of model training. To further mitigate one-to-one misalignments, ensemble learning can be exploited to reduce variability in pseudo-label selection in semi-supervised settings. Self-ensembling methods, such as temporal ensembling [15], aggregate the predictions from a single model across different training epochs. Although self-ensembling has been shown to enhance prediction consistency in semi-supervised settings, it can inadvertently introduce confirmation bias by imposing cross-iteration dependencies and exacerbating error propagation in the context of pseudo-labeling.
To address this issue, we propose a parallel ensembling approach that refines pseudo-labeled alignments by combining predictions from multiple OT-based models trained in parallel. By seeking consensus across multiple models, this approach reduces prediction variability, thus enhancing the overall quality and reliability of pseudo-labeled alignments.